OpenStudy (anonymous):

Determine if the lines 4x+y=6 and -3x+3y=-6 are parallel, perpendicular or neither.

OpenStudy (anonymous):

First, you want to put both equations in slope-intercept form (y=mx+b)

OpenStudy (texaschic101):

put them in y = mx + b form where m is the slope and b is the y intercept. 4x + y = 6 y = -4x + 6 slope = -4 and y intercept = 6 -3x + 3y = -6 3y = 3x - 6 y = x - 2 slope = 1 and y intercept = -2 parallel lines will have the same slope perpendicular lines will have negative reciprocal slopes...all that means is flip the slope and change the sign. Example : 2/1 and -1/2 are negative reciprocals and they are perpendicular. if the slopes and y intercepts are different, then they are neither parallel or perpendicular so what do you think the answer is ?